OpenStudy (anonymous):

Determine the standard form of the equation of the line that passes through (–7, 8) and (0, –8). (1 point)

OpenStudy (anonymous):

(-8-8)/(0+7)= -16/7 would be the slope then y=mx+b is the equation you start with y+8=-16/7(x-0) gives you y-8=-16/7x which gives you y=-16/7x+8 then standard form Ax+By=C so, 16/7x +y=-8